1. The average score on IQ tests is 100, with a standard deviation of 15. One of your clients scored 134. What percent of the population is likely to score higher?


Group of answer choices

98.81%


27.36%

1.16%

48.81%

2. You work in the admission department of a small liberal-arts college in the Midwest. Scores from a standardized test are used as the primary method of determination of eligibility. The mean of that standardized test is 609, and the standard deviation is 109. Only students in the top 25% are accepted. Can you accept someone with a score of 699?


Group of answer choices

Yes


No

3. The average sleep time in a college dorm is 423 minutes, with a standard deviation of 157 minutes. What are the lower and upper limits of sleep time for 95% of the dorm inhabitants?


Group of answer choices

Upper = 625.14, lower = 224.98


Upper = 555.13, lower = 225.36

Upper = 730.72, lower = 115.28

None of the above

4. The t statistic is derived from:

Group of answer choices

a. The mode


c. The z score

b. The median

d. Who the heck knows

5. The one sample t test is used for what kind of analyses?

Group of answer choices

To compare means for two groups of cases


To determine whether the mean and the standard deviation of a sample are correlated

To compare the means of two variables for a single group

To test whether the mean of a single variable differs from a specified constant

6. The paired-sample t test is used for what kind of analyses?

Group of answer choices

To test whether the mean of a single variable differs from a specified constant


To compare means for two groups of cases

To compare the means of two variables for a single group

To determine whether the mean and the standard deviation of a sample are correlated

7.

The independent samples t test is used for what kind of analyses?


Group of answer choices

To compare means for two groups of cases


To compare the means of two variables for a single group

To test whether the mean of a single variable differs from a specified constant

To determine whether the mean and the standard deviation of a sample are correlated

8. The sampling distribution of the mean is

Group of answer choices

The distribution of sample means over repeated samples


The population mean

The distribution of the population mean over many populations

The mean of the distribution of the sample.

9. If we fail to reject the null hypothesis in a t test we can conclude

Group of answer choices

That the null hypothesis is false


That we don’t have enough evidence to reject the null hypothesis.

That the null hypothesis is true

That the alternative hypothesis is false

Please help! I would greatly appreciate it.
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

Normal dist params

Mean = 100
Stdev = 15
X = 134
Lets use the normal params to standardize using the formula:
Z = (x-Mean)/(Stdev/sqrt(n))

P(higher than 134)
= P(X>134)

Standardizing the raw score X to Z:

= P(Z> (X-Mean)/Stdev)
= P(Z> (134-100)/15)
= P(Z>2.2667)
= 1-P(Z<=2.2667) [Lookup the Z table to convert a Z score to a cumulative prob]
= 1-0.9883
= 0.0116 or 1.16%

Answer: 1.16%